Question 1-4: Pretend that you have a computer with base 10 with 4 digits in the significand. This computer truncates (not round) after each arithmetic operation. For example, the sum of 24.57 + 128.3 = 152.87 becomes 152.8 Question 1 What is the result when 128.3, 3.163, 24.57, and 0.4825 are added in their original order? a. 156.5 b. 156.6 c. 156.52 d.156.5155 e. 156.3 f. 156.4 Question 2 What is the result when the same four numbers are added in ascending order (low to high)? a. 156.5 b. 156.6 c. 156.52 d. 156.5155 e. 156.3 f. 156.4 Question:3 What is the result when the same four numbers are added in descending order (high to low)?Please Explain why the answer is what it is and which of the three (original, ascending, or descending is the most accurate.
Answer #1

Ans for Q1)

Given numbers are 128.3,3.163,24.57 and 0.4825

If you add the values in original order

128.3 + 3.163 = 131.463 #Add first two values and trucate the result but not round off then value becomes 131.4

131.4 + 24.57 = 155.97 #Now add given add third value to above truncated result then value becomes 155.9

155.9 + 0.4825 = 156.3825 #Now again add forth value to above truncated result then value becomes 156.3

So ANS is 156.3

Ans for Q2)

Given numbers are 128.3,3.163,24.57 and 0.4825

Ascending order: 0.4825,3.163,24.57 and 128.3

0.4825 + 3.163 = 3.6455 #Add first two values and trucate the result but not round off then value becomes 3.6

3.6 + 24.57 = 28.17 #Now add given add third value to above truncated result then value becomes 28.1

28.1 + 128.3 = 156.4 #Now again add forth value to above truncated result then value becomes 156.4

So ANS is 156.4

Ans for Q3)

Given numbers are 128.3,3.163,24.57 and 0.4825

Descending order: 128.3,24.57,3.163 and 0.4825

128.3 + 24.57 = 152.87 #Add first two values and trucate the result but not round off then value becomes 152.8

152.8 + 3.163 = 155.963 #Now add given add third value to above truncated result then value becomes 155.9

155.9 + 0.4825 = 156.3825 #Now again add forth value to above truncated result then value becomes 156.3

So ANS is 156.3

Note: In your question which one of three is accurate---All three answers are correct as per three condions in above questions. If no condition given, you can add all four values in any order then ans becomes same in all cases.